# Eco-evolutionary constraints for the endemicity of rapidly evolving viruses

### v1

@@ Abstract
Antigenic escape constitutes the main mechanism allowing rapidly evolving viruses to achieve endemicity. Beyond granting immune escape, empirical evidence also suggests that mutations of viruses might increase their inter-host transmissibility. While both mechanisms are well-studied individually, their combined effects on viral endemicity remain to be explored. Here we propose a minimal eco-evolutionary framework to simulate epidemic outbreaks generated by pathogens evolving both their transmissibility and immune escape. Our findings uncover a very rich phenomenology arising from the complex interplay between both evolutionary pathways and the underlying contagion dynamics. We first show that contagions at the population level constrain the effective evolution of the virus, accelerating the increase in transmissibility in the first epidemic wave while favoring antigenic variation in the transition to the endemic phase. Our results also reveal that accounting for both evolutionary pathways changes the features of the viruses more prone to become endemic. While chances for endemicity increase with infectiousness of the wild-type variant for viruses not evolving their transmissibility, a non-monotonic behavior is observed when the latter mechanism is included, favoring less transmissible viruses and impairing those ones with intermediate infectiousness.
@@ Introduction
Yet eliciting an immune response in their hosts, some rapidly evolving viruses persist in real populations due to the antigenic variation arising from mutations in antibody-binding regions [8, 9, 10, 11]. This evolutionary mechanism facilitates reinfection events by antigenically distant variants that evade pre-existing immunity within the population. The relation between reinfection events and antigenic distances has been incorporated into theoretical models, assuming that variants can be embedded as nodes in genotype networks [12, 13, 14] or in discrete [15, 16, 17, 18, 17] or continuous [19, 20] low-dimensional representations of the antigenic space. Rouzine et al. [19] show that the interplay between immune presure and virus evolution yields travelling fitness waves in 1D antigenic spaces, sustaining enough viral diversity for the virus to persist in the population. Morever, antigenic spaces with higher dimensionality can also harbor the speciation of viruses, splitting at some point into different sublineages which evolve independently [20, 18].
@@
Pathogens’ evolution also shapes other non-antigenic traits such as their inter-host transmissibility [21, 22] or their virulence [23, 24], leading to major changes in their associated epidemic trajectories [25, 26, 27] and their impact on public health [28, 29]. For instance, recent genomic surveillance data reveals that SARS-CoV-2 virus has undergone both evolution in its transmissibility and antigenic escape [30, 31]. Motivated by this empirical evidence, several theoretical works have proven that the underlying eco-evolutionary dynamics for these viruses intertwine both evolutionary pathways. Namely, non-antigenic traits can evolve towards maximizing either the basic reproduction number ${\mathcal{R}_{0}}$ or the speed of the antigenic fitness wave at equilibrium as a function of the immune pressure existing in the population [32, 33].
@@
As the SARS-CoV-2 virus transitions from an epidemic to an endemic phase [34], there is increasing interest in determining the eco-evolutionary constraints that enable this transition in real populations [35, 36]. To the best of our knowledge, understanding how evolution in both antigenic and non-antigenic traits determines the endemicity of viruses remains an open question. To fill this gap, here we propose a minimal eco-evolutionary framework, extending the classical Susceptible-Infected-Recovered (SIR) model to integrate the evolution of pathogens in both their transmissibility and their immune escape. In absence of the former evolutionary pathway, we show that the probability of a virus becoming endemic depends on its antigenic escape and always increases with the infectiousness of the wild-type variant. In contrast, our stochastic simulations reveal that evolution in non-antigenic traits significantly alter this picture, resulting in a non-monotonic behavior where viruses with low infectiousness display greater endemicity than others with intermediate infectiousness. We derive a heuristic equation capturing this behavior as a result of the trade-off between the minimum infectiousness needed to sustain an endemic scenario and the characteristic evolutionary to reach such infectiousness. Finally, we discuss the practical implications of our findings.
@@ Results
.1 Model for the coevolution of transmissibility and antigenic escape
Fig. 1 sketches the main features of the minimal eco-evolutionary framework here introduced. We assume a constant and well-mixed population of $N$ individuals making $k$ contacts per time step. As stated above, from an epidemiological point of view, our framework is a modified version of the SIR model to harbor reinfection events for those variants escaping the immune response mounted in the population. In this sense, each virus $j$ circulating across the population is characterized by two variables: its transmissibility ${\lambda_{j}}$ and its position in the antigenic space $x_{j}$. For the sake of simplicity, we consider a 1D antigenic space, aligning with previous theoretical frameworks [33, 19].
Fig. 1A illustrates all the processes changing the epidemiological state of individuals in our model. Namely, each infected individual $j$ recover at a rate $\mu$, preserving the antigenic position $x_{j}$ of the variant for which they developed their immune response. Regarding contagions, we assume that any susceptible individual in contact with an infected individual $j$ contracts the pathogen at a rate $\lambda_{j}$. Conversely, the rate of reinfection events $\lambda^{\prime}_{ij}$ is different for each recovered individual $i$, as it should account for the antigenic distance between its associated variant and that of the infected agent. Following [15, 19], for each pair $(i,j)$ formed by a recovered individual $i$ interacting with an infected agent $j$, we assume that the contagion rate is given by:
@@
The latter choice imposes that reinfection events only occurs when $x_{j}>x_{i}$. Such assumption gives rise a travelling wave of variants moving with a positive speed in the 1D antigenic space [19]. Hence, our model assumes that natural immunity against past variants does not wane over time. For all contagions, the agent contracting the pathogen inherits the variant of the infected individual.
Figure 1: Schematic of the eco-evolutionary model here introduced. A) Modified SIR model to account for the reinfection events driven by the immune escape of the virus. Each infected agent $j$ recovers at a rate $\mu$, keeping their associated antigenic position $x?j$. Primary infections caused by an infected agent $j$ occur at a rate $\lambda_{j}$, whereas the contagion rate for a recovered individual $i$ in contact with that infected individual is $\lambda^{\prime}_{ij}=\Theta(x_{j}-x_{i})\lambda_{j}\left(1-e^{-(x_{j}-x_{i})}\right)$, where $\Theta(x)$ represents the Heavyside function. Note that the latter increases with the antigenic distance between variants and that reinfection is only possible when $x_{j}>x_{i}$, thus assuming that natural immunity against a variant does not wane over time. B) Evolution of the virus transmissibility and antigenic position in infected individuals, assuming that changes in trait $m$ ($m\in\left\{x,\lambda\right\}$) follows $\Delta_{m}\sim\mathcal{N}(0,D^{2}_{m})$, with $D^{2}_{m}$ determining its speed of evolution.
@@
Fig. 1B shows the evolutionary processes changing the epidemiological parameters of the circulating viral strains. We assume that mutations just occur in infected individuals, shaping both the transmissibility of the virus and its antigenic position by a factor $\Delta\lambda$ and $\Delta x$ respectively. Inspired by previous works [25, 19, 32], we assume that changes in transmissibility (antigenic position) are drawn from a zero-mean Gaussian distribution whose variance $D^{2}_{\lambda}$ ($D^{2}_{x}$) determines the speed of evolution. Regarding recovered individuals, we assume their antigenic variant to remain constant over time until they get reinfected by another circulating strain. A complete description of the stochastic simulations performed to obtain the epidemic trajectories following this model is provided in the Materials and Methods section.
.2 Endemicity of viruses without transmissibility evolution
To understand the transition from the epidemic to the endemic phase, let us first neglect the evolution of the virus in the transmissibility space by setting $D_{\lambda}=0$ and simulate different epidemic outbreaks varying the speed of evolution in the antigenic space $D_{x}$. Unless otherwise stated, throughout the manuscript we fix the recovery rate of the disease to $\mu=1.0/7$ days-1 and consider epidemic outbreaks spreading across populations of $N=10^{4}$ individuals, each one making $k=10$ contacts per day.
Fig. 2A depicts the three different epidemic regimes which can be found in the model for a pathogen whose wild-type variant has a basic reproduction number $\mathcal{R}^{wt}_{0}=\lambda k/\mu=3$. Recall that the basic reproduction number quantifies the number of contagions triggered by one infected individual in a healthy population. Therefore, when $\mathcal{R}^{wt}_{0}>1$, the pathogen always trigger a first epidemic wave in the population. Nonetheless, the long-term behavior of the epidemic trajectories does depend on the immune escape mechanism, parametrized with $D_{x}$. In absence of immune escape, i.e. $D_{x}=0$, the model reduces to the standard SIR model characterized by a single epidemic wave followed by the extinction of the infected population. Conversely, for $D_{x}=0.03$, a damped oscillatory behavior towards an endemic equilibrium appears for all epidemic trajectories. In between these two extremes ($D_{x}=0.015$), richer epidemic dynamics are generated by the model, as the first epidemic wave is typically followed by long periods of very low epidemic incidence. Such periods act as epidemic bottlenecks, leading to the extinction of some outbreaks due to stochastic fluctuations. Others, however, are able to persist in the population as a result of the enlargement of the distribution of variants in the antigenic space, as shown in Fig. 2B. Comparing the endemic realizations, higher $D_{x}$ values lead to endemic scenarios with a more acute prevalence of the diseases, as reported in [19].
@@
Figure 2: Epidemic trajectories for viruses not evolving their transmissibility. A) Time dependence of the fraction of population in the infected state $\rho_{I}$ as a function of the speed of evolution in the antigenic space $D_{x}$ (color code). B): Time evolution of the variance of the distribution of variants in the antigenic space $\sigma_{x}^{2}$ as a function of $D_{x}$ (color code). In both panels, thick solid (dashed) lines shows the values for those viruses becoming (not becoming) endemic in the population obtained by averaging 200 epidemic outbreaks whereas thin lines represent a sample of 20 individual trajectories in both cases. In these panels, the basic reproduction number of the wild-type variant of the pathogen is set to $\mathcal{R}^{wt}_{0}$=3. C): Fraction of epidemic outbreaks surviving in the population after $t=1000$ days as a function of ${\mathcal{R}^{wt}_{0}}$ and $D_{x}$ (color code). The results shown in this panel have been obtained by simulating $1000$ epidemic outbreaks for each pair of ($\mathcal{R}^{wt}_{0}$,$D_{x}$) values. In all panels, we assume $N=10^{4}$ individuals, $I_{0}=10$ initially infected agents, and we set the recovery rate to $\mu=1/7$ days-1 and the number of daily contacts to $k=10$ interactions.
@@
To further deepen into the features shaping virus’ endemicity, Fig. 2C shows the fraction of epidemic outbreaks surviving in the population after $t=1000$ days as a function of both ${\mathcal{R}_{0}}^{wt}$ and $D_{x}$. There, the three aforementioned epidemic regimes become more evident. Moreover, virus endemicity increases monotonically with the infectiousness of the wild-type variant regardless of the speed of evolution in the antigenic space $D_{x}$. Likewise, for a given $\mathcal{R}_{0}^{wt}$ value, accelerating the evolution in the antigenic space, i.e. increasing $D_{x}$, also enhances virus endemicity, making immune escape more likely.
.3 Endemicity of viruses with transmissibility evolution
@@
So far, we have neglected the evolution of virus transmissibility, retrieving the three dynamical regimes obtained in the literature for the SIR model with immune escape and the expected monotonic behavior of endemicity with virus infectiousness. Hereafter, we include such evolutionary pathway, questioning how the joint evolution of both antigenic and non-antigenic traits alters this picture. We first fix the evolution speed of the different traits by setting $D_{x}=0.015$ and $D_{\lambda}=3\cdot 10^{-4}$ and study the epidemic trajectories for three different pathogens differing in the basic reproduction number of their wild type variants ${\cal R}_{0}^{wt}$.
@@
Fig. 3A shows that evolution in transmissibility allows viruses with low initial infectiousness, e.g. ${\cal R}_{0}^{wt}=1.25$ or ${\cal R}_{0}^{wt}=2$, to become endemic in the population. Note that this result comes in stark contrast with those reported in Fig. 2C, as no endemic realization is observed for this combination of parameters when just accounting for the evolution in antigenic traits. Regarding the shape of the temporal trajectories, the early stages of the outbreak qualitatively resemble the ones shown in Fig. 2A with the presence of a first prominent epidemic wave followed by an epidemic bottleneck. Conversely, the late stages of the outbreak are not characterized by a fixed prevalence of the disease as in the former case. Interestingly, regardless of the infectiousness of the wild-type variant, the epidemic prevalence seems to grow at a (roughly) constant pace on average once the virus has entered the endemic phase.

### v2

@@ Abstract
Antigenic escape constitutes the main mechanism allowing rapidly evolving viruses to achieve endemicity. Beyond granting immune escape, empirical evidence also suggests that mutations of viruses might increase their inter-host infectiousness. While both mechanisms are well-studied individually, their combined effects on viral endemicity remain to be explored. Here we propose a minimal eco-evolutionary framework to simulate epidemic outbreaks generated by pathogens evolving both their infectiousness and immune escape. Our results reveal that the main driver of viral evolution shifts over time: from intrinsic selection for infectiousness at early stages of the outbreak to antigenic diversification in the transition to the endemic phase. We find that the evolution in both traits during the first epidemic wave plays a critical role in determining long-term viral persistence. Evolution in infectiousness enhances the endemicity of viruses, especially in viruses with lower baseline infectiousness due to the longer duration of their first epidemic wave. Likewise, control policies flattening epidemic curves might increase viral endemicity as a result of the greater antigenic diversity generated in the prolonged epidemic waves. Our results thus prove that the long-term behavior of epidemic trajectories hinges on the complex interplay between both evolutionary pathways and the underlying contagion dynamics.
@@ Introduction
Yet eliciting an immune response in their hosts, some rapidly evolving viruses persist in real populations due to the antigenic variation arising from mutations in antibody-binding regions [8, 9, 10, 11]. This evolutionary mechanism facilitates reinfection events by antigenically distant variants that evade pre-existing immunity within the population. The relation between reinfection events and antigenic distances has been incorporated into theoretical models, assuming that variants can be embedded as nodes in genotype networks [12, 13, 14] or in discrete [15, 16, 17, 18] or continuous [19, 20] low-dimensional representations of the antigenic space. Rouzine et al. [19] show that the interplay between immune presure and virus evolution yields travelling fitness waves in 1D antigenic spaces, sustaining enough viral diversity for the virus to persist in the population. Moreover, antigenic spaces with higher dimensionality can also harbor the speciation of viruses, splitting at some point into different sublineages which evolve independently [20, 18].
@@
Pathogens’ evolution also shapes other non-antigenic traits such as their inter-host transmissibility [21, 22] or their virulence [23, 24], leading to major changes in their associated epidemic trajectories [25, 26, 27] and their impact on public health [28, 29]. For instance, recent genomic surveillance data reveals that SARS-CoV-2 virus has undergone both evolution in its infectiousness and antigenic escape [30, 31]. Motivated by this empirical evidence, several theoretical works have proven that the underlying eco-evolutionary dynamics for these viruses intertwine both evolutionary pathways. Namely, non-antigenic traits can evolve towards maximizing either the basic reproduction number $\mathcal{R}_{0}$} or the speed of the antigenic fitness wave at equilibrium as a function of the immune pressure existing in the population [32, 33].
Figure 1: Schematic of the eco-evolutionary model here introduced. A) Modified Susceptible-Infected-Recovered (SIR) model to account for the reinfection events driven by the immune escape of the virus. Each agent $j$, when infected, recovers at a rate $\mu$, keeping their associated antigenic position $x_{j}$. Primary infections caused by an infected agent $j$ occur at a rate $\lambda_{j}$, whereas the contagion rate for a recovered individual $i$ in contact with that infected individual is $\lambda^{\prime}_{ij}=\Theta(x_{j}-x_{i})\lambda_{j}\left(1-e^{-(x_{j}-x_{i})}\right)$, where $\Theta(x)$ represents the Heaviside function. Note that the latter increases with the antigenic distance between variants and that reinfection is only possible when $x_{j}>x_{i}$, thus assuming that natural immunity against a variant does not wane over time. B) Evolutionary processes of the model. On the one hand, recovered individuals keep their antigenic position. On the other hand, infected individuals evolve both the infectiousness of the virus and their antigenic position, assuming that changes in trait $m$ ($m\in\left\{x,\lambda\right\}$) are drawn from normal distributions, i.e. $\Delta_{m}\sim\mathcal{N}(0,D^{2}_{m})$, with $D^{2}_{m}$ determining its speed of evolution.
@@
As the SARS-CoV-2 virus transitions from an epidemic to an endemic phase [34], there is increasing interest in determining the eco-evolutionary constraints that enable this transition in real populations [35, 36]. To the best of our knowledge, understanding how evolution in both antigenic and non-antigenic traits determines the endemicity of viruses remains an open question. To fill this gap, here we propose a minimal eco-evolutionary framework, extending the classical Susceptible-Infected-Recovered (SIR) model to integrate the evolution of pathogens in both their infectiousness and their immune escape. In absence of the former evolutionary pathway, we show that the probability of a virus becoming endemic depends on its antigenic escape and always increases with the infectiousness of the wild-type variant. In contrast, our stochastic simulations reveal that evolution in non-antigenic traits alters this picture, resulting in a non-monotonic behavior where viruses with low infectiousness display greater endemicity than others with intermediate infectiousness. We derive a heuristic equation capturing this behavior as a result of the trade-off between the minimum infectiousness needed to sustain an endemic scenario and the characteristic evolutionary time scale to reach such infectiousness. To control this evolutionary time scale, we implement control policies flattening epidemic curves and report that extending the first epidemic wave leads to an increase in viral endemicity. To round off the manuscript, we also show discuss the practical implications of our findings.
@@ Results
Model for the coevolution of infectiousness and antigenic escape
Fig. 1 sketches the main features of the minimal eco-evolutionary framework here introduced. We assume a constant and well-mixed population of $N$ individuals, where each individual draws $k$ randomly chosen contacts each time step. Our framework is a modified version of the SIR model to harbor reinfection events for those variants escaping the immune response mounted in the population. In this sense, each individual $j$ is characterized by its position in the antigenic space $x_{j}$, denoting the immune response developed against the virus. For the sake of simplicity, we consider a 1D antigenic space, aligning with previous theoretical frameworks [33, 19]. In addition, we assume that if the individual $j$ is infected, they carry a single variant of the virus, characterized by its infectiousness $\lambda_{j}$.
Fig. 1A illustrates all the processes changing the epidemiological state of individuals in our model. Namely, each infected individual $j$ recovers at a rate $\mu$, preserving the antigenic position $x_{j}$ of the variant for which they developed their immune response. Regarding contagions, any susceptible individual in contact with an infected individual $j$ contracts the pathogen at a rate $\lambda_{j}$. Conversely, the rate of reinfection events $\lambda^{\prime}_{ij}$ is different for each recovered individual $i$, as it should account for the distance between its associated antigenic position and that of the infected agent.
Following [15, 19, 32], for each pair $(i,j)$ formed by a recovered individual $i$ interacting with an infected agent $j$, we assume that the contagion rate is given by:
@@
Consequently, reinfection events only occur when $x_{j}>x_{i}$. This assumption is needed to avoid reinfections of variants already contracted in the past and account for the immune memory host. For all contagions, the agent contracting the pathogen inherits the variant of the infected individual.
@@
Fig. 1B shows the evolutionary processes changing the epidemiological parameters of the circulating viral strains. We assume that mutations just occur in infected individuals. To reflect the impact of these mutations, for each time step, we update the antigenic position of each infected individual $j$, $x_{j}$, by a factor $\Delta x$. Inspired by previous works [25, 19, 32], we assume that changes in antigenic position are drawn from a zero-mean Gaussian distribution whose variance $D^{2}_{x}$ determines the speed of evolution. Regarding recovered individuals, we assume their antigenic position to remain constant over time until they get reinfected by another circulating strain. Note that negative changes in antigenic position, yet occurring in mutations inside hosts, are not selected at the population level and the overall behavior corresponds to a travelling wave of strains moving at a positive speed in the 1D antigenic space [19]. Moreover, we assume that changes in infectiousness are drawn from another uncorrelated zero-mean Gaussian distribution whose variance is denoted by $D^{2}_{\lambda}$. A complete description of the stochastic simulations performed to obtain the epidemic trajectories shown in the manuscript is given in Appendix A.
Endemicity of viruses without evolution in infectiousness
@@
Figure 2: Epidemic trajectories for viruses not evolving their infectiousness. A) Time dependence of the fraction of population in the infected state $\rho_{I}$. B): Time evolution of the variance of the distribution of variants in the antigenic space $\sigma_{x}^{2}$. In both panels, thick solid (dashed) lines shows the values for those viruses becoming (not becoming) endemic in the population obtained by averaging 200 epidemic outbreaks whereas thin lines represent a sample of 20 individual trajectories in both cases. In these panels, the basic reproduction number of the wild-type variant of the pathogen $\mathcal{R}_{0}^{wt}$} is set to $\mathcal{R}^{wt}_{0}$=3. C): Fraction of epidemic outbreaks surviving in the population after $t=1000$ days $f_{endemic}$ as a function of $\mathcal{R}^{wt}_{0}$}. The results shown in this panel have been obtained by simulating 10001000 epidemic outbreaks for each pair of ($\mathcal{R}^{wt}_{0}$$D_{x}$) values. In all panels, line color represents the value of speed of evolution in antigenic position $D_{x}$. For the simulations, in all panels we assume $N=10^{4}$ individuals, $I_{0}=10$ initially infected agents, and we set the recovery rate to $\mu=1/7$ days-1 and the number of daily contacts to $k=10$ interactions.
To understand the transition from the epidemic to the endemic phase, let us first neglect the evolution of the virus in infectiousness by setting $D_{\lambda}=0$ and simulate different epidemic outbreaks varying the speed of evolution in the antigenic space $D_{x}$. Unless otherwise stated, throughout the manuscript we fix the recovery rate of the disease to $\mu=1.0/7$ days-1 and consider epidemic outbreaks spreading across populations of $N=10^{4}$ individuals, each one making $k=10$ contacts per day. For all the simulations, we assume an initial prevalence of the disease $\rho_{I}(0)=0.001$ on average and consider that all the individuals initially infected share the same wild-type strain $j$, characterized by $x_{j}=0$ and $\lambda_{j}={\cal R}_{0}^{wt}\mu/k$, where ${\cal R}_{0}^{wt}$ denotes its basic reproduction number.
Fig. 2A depicts the three different epidemic regimes which can be found in the model for a pathogen with $\mathcal{R}^{wt}_{0}=3$. Recall that the basic reproduction number quantifies the number of contagions triggered by one infected individual in a healthy population. Therefore, when $\mathcal{R}^{wt}_{0}>1$, the pathogen always produces a first epidemic wave in the population. Nonetheless, the long-term behavior of the epidemic trajectories does depend on the immune escape mechanism, parametrized with $D_{x}$. In absence of immune escape, i.e. $D_{x}=0$, the model reduces to the standard SIR model characterized by a single epidemic wave followed by the extinction of the infected population. Conversely, for $D_{x}=0.03$, a damped oscillatory behavior towards an endemic equilibrium appears for all epidemic trajectories. In between these two extremes ($D_{x}=0.015$), richer epidemic dynamics are generated by the model, as the first epidemic wave is typically followed by long periods of very low epidemic incidence. Such periods act as epidemic bottlenecks, leading to the extinction of some outbreaks due to stochastic fluctuations. Others, however, are able to persist in the population as a result of the enlargement of the distribution of variants in the antigenic space, as shown in Fig. 2B. Comparing the endemic realizations, higher $D_{x}$ values lead to endemic scenarios with a more acute disease prevalence, as reported in [19].
@@
Fig. 2C shows the fraction of epidemic outbreaks surviving in the population after $t=1000$ days, denoted in what follows by $f_{endemic}$, as a function of both ${\mathcal{R}}_{0}^{wt}$ and $D_{x}$. There, the three aforementioned epidemic regimes become more evident. Moreover, virus endemicity increases monotonically with the basic reproduction number of the wild-type variant regardless of the speed of evolution in the antigenic space $D_{x}$. Likewise, for a given $\mathcal{R}_{0}^{wt}$ value, accelerating the evolution in the antigenic space, i.e. increasing $D_{x}$, also enhances virus endemicity, making immune escape more likely.
Endemicity of viruses with evolution in infectiousness
@@
So far, we have neglected the evolution of virus infectiousness, retrieving the three dynamical regimes obtained in the literature for the SIR model with immune escape and the expected monotonic behavior of endemicity with virus infectiousness. Hereafter, we include such evolutionary pathway and address how the joint evolution of both antigenic and non-antigenic traits alters this picture. We first fix the evolution speed of the different traits by setting $D_{x}=0.015$ and $D_{\lambda}=3\cdot 10^{-4}$ and study the epidemic trajectories for three different pathogens differing in the basic reproduction number of their wild type variants ${\cal R}_{0}^{wt}$.
@@
Fig. 3A shows that evolution in infectiousness allows mildy infectious viruses, e.g. ${\cal R}_{0}^{wt}=1.25$ or ${\cal R}_{0}^{wt}=2$, to become endemic in the population. Note that this result comes in stark contrast with those reported in Fig. 2C, as no endemic realization is observed for this combination of parameters when just accounting for the evolution in antigenic traits. Regarding the shape of the temporal trajectories, the early stages of the outbreak qualitatively resemble the ones shown in Fig. 2A with the presence of a first prominent epidemic wave followed by an epidemic bottleneck. Conversely, the late stages of the outbreak are not characterized by a fixed prevalence of the disease as in the former case. Interestingly, regardless of the infectiousness of the wild-type variant, the epidemic prevalence seems to grow at a (roughly) constant pace on average once the virus has entered the endemic phase.